OpenStudy (anonymous):

The path that a ball follows when thrown into the air from a height of 5 feet is given by the formula h=-0.1d^2 + 1.5d + 5, where h is the height of the ball and d is the time in second. How long will it take the ball to reach 10 feet? Please help it needs to be solved algebraically.

OpenStudy (anonymous):

\[10=-.1d^2+1.5d+5\] first get all everything to one side \[10-10=-.1d^2+1.5d+5-10\] \[0=-.1d^2+1.5d-5\]

OpenStudy (anonymous):

now use the quadratic formula which you can use when you have \[0=ax^2+bx+c\]

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

does this hep?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

btw you need to take the lower value, because it will reach 10 feet on it's way down too!

OpenStudy (anonymous):

oh okay! thanks!

OpenStudy (anonymous):

i have a question what do ou do after the quadratic formula?

OpenStudy (anonymous):

you should get 2 values for d

OpenStudy (anonymous):

did you find them?

OpenStudy (anonymous):

after you put everything in, you should be able to evaluate it.... so since there is a plus or minus after negative b, you'll get two answers one \[\frac{-b+\sqrt{b^2-4ac}}{2a}\] and \[\frac{-b-\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

so does x=10 and x=5?

OpenStudy (anonymous):

not quite sure let me check ... http://www.wolframalpha.com/input/?i=-.1x%5E2%2B1.5x-5%3D0 yep you're correct so x=d=t so thats your time

OpenStudy (anonymous):

so 5 seconds after you toss it, it gets to 10 feet then goes higher. Another 5 seconds later the ball fall back to 10 feet.

OpenStudy (anonymous):